# Draft: May 24, 2025

### v1

@@ 6 Suggestions for Users and Developers
To develop the library, consider to follow the procedure described above for a user with the only difference that, in your case, one or more thermodynamic compartments of $\mathcal{N}_{\beta}$ are not available in CiRL, and hence, you need to implement them yourself. In the case you need a robotic compartment, the best way to proceed is to leverage and modify as needed an existing physics engine as we did in FetchWasteSorting and Reacher (Table 1). In contrast, the literature lacks of RL physics engines for non-robotic systems such as those implemented in CiRL, i.e., the incinerator, the transportation truck, and the microalgae cultivations. In the non-robotic case, the advice is to formulate the compartment model in the state-space form similarly to CiRL since it is a standard format used by the automatic control community useful for systems analysis and controller designs. The ordinary differential equations in the state-space form have to be derived from mass balances and/or the laws of thermodynamics in order to yield a thermodynamic compartment of a TMN (Definition 2), and hence, to align with the CiRL framework. This is always possible thanks to the generality of thermodynamics (Haddad, 2017), although in some cases it can be particularly challenging. Then, the integration step size for numerical integration of the differential equations must be tuned as we did, for example, with TransportTruck and CO2MicroalgaeDroop (Table 1). The suitability of the integration step size can be verified by comparing the state trajectories produced by implementing the model as an SB3 environment (numerically solved via, for example, the Euler’s method) and the state trajectories given by the Python solver scipy.integrate.odeint() for the same initial conditions and the same control input (i.e., the action). This is what we did, for example, with TransportTruck and CO2MicroalgaeDroop. Once the state-space form is formulated and solved with a numerically-stable method, the reward function has to be defined in order to, directly or indirectly, increase the circularity $\lambda$ of $\mathcal{N}_{\beta}$. Finally, one or more SB3 algorithms must be called to train and test the RL controller optimizing the circularity-oriented reward. The newly implemented compartments could be added to CiRL to expand the library for future uses.
@@ 7 Limitations and Future Work
The current limitation of CiRL is that $\lambda(\mathcal{N};t)$ does not enter explicitly the reward functions of the environments, which would be the most natural way to optimize circularity via reinforcement learning. In contrast, the reward functions of the existing environments affect $\lambda(\mathcal{N};t)$ indirectly. This is because an explicit form of $\lambda(\mathcal{N};t)$ requires to preliminarily define a compartmental network $\mathcal{N}$, which is a network of environments/compartments. Controlling such a multi-environment scenario is clearly more complex than controlling separately the compartments as done here. Improving CiRL by including the control of multi-compartment sytems, and hence, to optimize $\lambda(\mathcal{N};t)$ directly, is the main future research direction. This first version of CiRL is a necessary step.
Another future work will be to target specific material flows, e.g., rare-Earth elements (European Commission, b) or critical raw materials (European Commission, 2023), and then, design networks $\mathcal{N}$ of compartments processing those materials whose circularity $\lambda(\mathcal{N};t)$ is optimized via DRL. This will result in extending the library of compartments available in CiRL.

### v2

@@ 6 Suggestions for Users and Developers
The aim of CiRL is to serve as a tool for decision makers interested in material flow circularity for a more sustainable management of natural resources. Specifically, a trained agent can be implemented in the compartmental network $\mathcal{N}$ and generate a sequence of actions. In practice, the actions produced by the agent are discrete-time values of material flows that can be replicated in real scenario. Clearly, the quality of the actions is as good as the realisticness of the RL environment that the agent has been trained on. At the current stage, our RL environments are provided as paradigmatic examples and do not model any real systems with fidelity; this has been necessary to introduce the concept of CiRL, and hence, to illustrate the use of state-of-the-art RL in the design of circular systems. Making these RL environments more realistic is essential to make CiRL useful for practical scenarios. A guide on how to use and develop CiRL is the scope of this section.
@@
To develop the library, one should consider to follow the procedure described above for a user with the only difference that, in this case, one or more thermodynamic compartments of $\mathcal{N}_{\beta}$ are not available in CiRL, and hence, one needs to implement them. In the case one needs a robotic compartment, the best way to proceed is to leverage and modify as needed an existing physics engine as we did in FetchWasteSorting and Reacher (Table 1). In contrast, the literature lacks of RL physics engines for non-robotic systems such as those implemented in CiRL, i.e., the incinerator, the transportation truck, and the microalgae cultivations. In the non-robotic case, the advice is to formulate the compartment model in the state-space form similarly to CiRL since it is a standard format used by the automatic control community useful for systems analysis and controller designs. The ordinary differential equations in the state-space form have to be derived from mass balances and/or the laws of thermodynamics in order to yield a thermodynamic compartment of a TMN (Definition 2), and hence, to align with the CiRL framework. This is always possible thanks to the generality of thermodynamics (haddad2017thermodynamics), although in some cases it can be particularly challenging. Then, the integration step size for numerical integration of the differential equations must be tuned as we did, for example, with TransportTruck and CO2MicroalgaeDroop (Table 1). The suitability of the integration step size can be verified by comparing the state trajectories produced by implementing the model as an SB3 environment (numerically solved via, for example, the Euler’s method) and the state trajectories given by the Python solver scipy.integrate.odeint() for the same initial conditions and the same control input (i.e., the action). This is what we did, for example, with TransportTruck and CO2MicroalgaeDroop. Once the state-space form is formulated and solved with a numerically-stable method, the reward function has to be defined in order to, directly or indirectly, increase the circularity $\lambda$ of $\mathcal{N}_{\beta}$. Finally, one or more SB3 algorithms must be called to train and test the RL controller optimizing the circularity-oriented reward. The newly implemented compartments could be added to CiRL to expand the library for future uses.
7 Applications and Impact
CiRL is a library developed as a support tool in the design of circular systems. The need of such systems can be considered strategic in the increasing competition between nations to secure access to critical raw materials necessary to make green, digital, and defense technologies (competition1; nygaard2023geopolitical). The traditional (i.e., linear) economy has also problems at the output stage since it generates large volumes of waste whose management is damaging the natural environment (egger2025evaluating; cottom2024local). Therefore, actions to increase the circularity of materials is gaining importance (macleod2024waste), even more considering the fact that the world population is expected to increase and reach approximately 10 billion in 2060 (owid-un-population-2024-revision).
The purpose of RL is to generate AI-driven actions so that the RL environment behaves as desired. By developing CiRL, which aims to optimize the systems circularity, we pave the way for generating AI-driven decisions to increase circularity. The state-of-the-art approach to decision making for circularity is material flow analysis (MFA), which is essentially a methodology based on the analysis of large data of material stocks and flows (cullen2022material; brunner2016handbook; EU-MFA; CSIRO-MFA). Hence, the traditional MFA approach consists of actions for circularity defined by humans formulated on the analysis of data. In constrast, CiRL leverages ordinary differential equations to define the models and, via RL, generates AI-driven actions for circularity. Thus, CiRL is a tool that decision makers can use in combination with MFA, especially considering the pace of improvement that AI systems have been showing in the last fifteen years.
The CiRL core is a general framework that can be applied to any sector, e.g., aerospace, healthcare, automotive, defense. Each sector involves the supply-recovery chain of several materials and products; thus, targeting a sector of interest requires to focus on the life-cycle stages, i.e., compartments, of the supply-recovery chains of its materials and products, and hence, to develop the CiRL environments modeling the material life-cycles of the target sector.
The current size of the RL environments in CiRL is limited, and hence, not suitable for realistic decision making. However, CiRL is modular, and hence, it is meant to be scaled-up to target specific real-world scenarios. Given that CiRL is the first library to sit between AI-driven decision making and circular economy, and given that leading European organizations are seeking circular designs solutions based on digital technologies (DiCE-Lab; DICENetwork+), the expected impact of CiRL is relatively high.
@@ 8 Limitations and Future Work
The current limitation of CiRL is that $\lambda(\mathcal{N};t)$ does not enter explicitly the reward functions of the environments, which would be the most natural way to optimize circularity via reinforcement learning. In contrast, the reward functions of the existing environments affect $\lambda(\mathcal{N};t)$ indirectly. Improving CiRL by including the control of multi-compartment systems optimizing $\lambda(\mathcal{N};t)$ directly is one of the main future research directions along with testing the trained agents in generating actions $u$ when interacting with the environment in a simulation loop. This is intended to be the first version of CiRL.
Other directions of future work are possible:
  • 1.
    Compare the performance of these RL-based controllers with techniques based on control theory, e.g., the linear quadratic regulator (LQR) (zocco2025carbonet).
  • 2.
    Develop inverse reinforcement learning systems to infer better reward functions for real-world applications (lanzaro2025evaluating; feng2025deep).
  • 3.
    Develop multi-agent RL agents to control multiple compartments simultaneously; for multi-agent RL, there might be libraries more suitable than SB3, e.g., MARLlib (hu2023marllib).
  • 4.
    Integrate explainability tools into CiRL to increase the interpretability of the RL agent actions; this is important in the case the AI-generated actions are considered for use in real case studies.
  • 5.
    Create and control compartmental networks targeting rare-Earth elements (Rare-EU) or critical raw materials (CRM-EU) similarly to what we did with CopperNet.
9 Conclusion
In this paper, we proposed CiRL, a library of reinforcement-learning environments that, by leveraging the learning algorithms in Stable-Baselines3, can provide AI-driven actions for optimizing the circularity of supply-recovery chains to be combined with human-driven decisions derived from material flow analysis (MFA) studies. CiRL currently consists of eight environments on which nine RL algorithms were tested. The results demonstrate the usability and scalability of CiRL. The main future work is to enhance the environment fidelity for specific case studies and to design reward functions optimizing circularity explicitly.